OpenStudy (anonymous):

cos^2x=(1/4) solve for x

OpenStudy (mathmate):

Hints: \( cos^2(x)=\frac{1}{4} \) then \( cos(x)=\pm \frac{1}{2} \) So you should expect 4 zeroes between 0 and 2\(\pi\).

OpenStudy (anonymous):

thank you!

OpenStudy (mathmate):

You're welcome!